How many grams of benzoic acid (C6H5COOH) are contained in a 125.0 mL solution at 0.26 M concentration?


1
Expert's answer
2021-05-19T02:27:43-0400

We know that


Molarity = "\dfrac{w*1000}{m*V}" .........................................Equation(1)


where, w = weight of the solute in the solution.

m = molar mass of the solute.

V = volume of the solution ( in mL).


It is given that, Molarity = 0.26 M

V = 125.0 mL

m = 122.12 gm


From Equation(1) we have,


0.26 = "\dfrac{w*1000}{122.12*125.0}"


w = "\dfrac{0.26*122.12*125.0}{1000}"

w = 3.9689 gm



3.9689 grams of benzoic acid (C6H5COOH) are contained in a 125.0 mL solution at 0.26 M concentration.